# Visual Sparse Steering: Improving Zero-shot Image Classification with Sparsity Guided Steering Vectors

### v1

@@ Appendix B Sensitivity to Sparse Amplification ($\gamma$) and Steering Magnitude ($\lambda$)
Figure 3 shows that all tested combinations of $(\lambda,\gamma)$ outperform the zero-shot baseline, though to varying degrees. Each dataset exhibits a diagonal band of near-optimal settings where $\lambda\cdot\gamma\in[2,3]$ tends to yield peak accuracy. For example, CIFAR-100 peaks at $\lambda^{*}=2.1$ and $\gamma^{*}=1.5$. Beyond this band, increasing $\lambda$ or $\gamma$ causes performance to degrade likely due to over-amplification of sparse codes and embedding distortion. The consistent contour patterns across datasets suggest that VS2 is robust to moderate variation in its hyperparameters and that good settings generalize well across domains.
@@ Appendix C Training top-$k$ Sparse Autoencoders
We follow CLIP [18] with a ViT-B/32 and ViT-B/16 vision backbones, intercepting the output of each encoder layer for the $CLS$ token. Specifically, we train top-$k$ SAEs on the $CLS$ embeddings for each chosen layer. We use $k=64$ and $k=256$ as the maximum active features within the latent space for ViT-B/32 and ViT-B/16 respectively, and we set a “dead feature” threshold of 100 i.e., any feature seldom activated is pruned. We also use an expansion factor of 4 relative to the input embedding dimension, resulting in 3,072 latent units. Training largely follows [31] and uses  [53], with a linear learning-rate schedule and warmup from $5\times 10^{-4}$.
@@
where $\mathbf{X}\in\mathbb{R}^{B\times d}$ is a batch of CLS embeddings, $\hat{\mathbf{X}}$ denotes their SAE reconstructions, and $\bar{\mathbf{X}}$ is the batch mean (so the denominator equals the total variance). FVU is the complement of the coefficient of determination ($1-R^{2}$); an FVU of 00 indicates perfect reconstruction, while an FVU of $1$ corresponds to predicting only the mean. In Table 4, we report FVU results for ViT-B/32 with $k=64$ and ViT-B/16 with $k=256$ across all three datasets.
@@ Appendix D Sensitivity in Retrieval-Augmented Generation (RAG)
When an external image corpus is available, VS2 can be extended using a Retrieval-Augmented Generation (RAG) pipeline. We use DINOv2 [50] to retrieve top-$k$ images most similar to the input query and compute an enhanced embedding by combining the query with the retrieved set:
@@
The parameter $\alpha\in[0,1]$ controls the trade-off between using the original query and the retrieved set.
We sweep over values of $\alpha$ and $k$ to assess their impact on zero-shot classification performance. Figure 4 shows results for CIFAR-100, and Tiny-ImageNet.
Findings.
CIFAR-100 and Tiny-ImageNet display a similar trend: larger $\alpha$ (i.e., more reliance on the query) typically degrades performance. On Tiny-ImageNet, setting $\alpha=0$, completely ignoring the input query and relying purely on retrieved features, yields the best result. This suggests that the retrieval set carries strong distributional signal in cluttered datasets. Across all datasets, large $k$ eventually hurts performance, confirming that RAG benefits from focused rather than broad context.
The trade-off parameter $\alpha$ and the retrieval depth $k$ play dataset-dependent roles in RAG-enhanced pipelines. For fine-grained domains, fewer, high-confidence neighbors and a low $\alpha$ work best; for noisy domains, more reliance on the retrieval pool is beneficial.
@@
Appendix E Are exemplar-derived directions really distinct?
@@
A desirable property of class–specific steering vectors is orthogonality: pushing an embedding toward class A should not simultaneously raise its score for class B. Using the oracle prototypes as described in Section 4.2, we compute for every CIFAR-100 class a prototype steering vector and measure the pair-wise cosine similarity at layer 11 of the ViT-B/32 encoder. Most pairs have low similarity ($\!\text{mean}\!=\!0.23$), yet a non-negligible tail reveals strongly overlapping directions. Table 5 lists the ten highest-overlap pairs:
@@
Appendix F Ablation Study of Expansion factor and top-$k$
SAE configuration Top-1 $\uparrow$ Top-5 $\uparrow$
4 $\times$, $k$=128 64.61 87.95
8 $\times$, $k$=128 64.56 87.76
4 $\times$, $k$=64 64.54 87.79
16 $\times$, $k$=64 64.54 87.78
8 $\times$, $k$=64 64.52 87.96
10 $\times$, $k$=128 64.43 87.81
16 $\times$, $k$=512 64.42 87.71
8 $\times$, $k$=512 64.40 87.91
4 $\times$, $k$=256 64.34 87.62
8 $\times$, $k$=256 64.29 87.80
16 $\times$, $k$=256 64.28 87.75
4 $\times$, $k$=512 64.12 87.87
16 $\times$, $k$=128 64.10 87.79
@@ Appendix F Ablation Study of Expansion factor and top-$k$
SAE cfg.
16×/64
16×/512
16×/128
16×/256
8×/128
8×/512
8×/256
4×/256
4×/512
4×/128
10×/128
4×/64
8×/64
16×/64 1.000 0.497 0.577 0.705 0.726 0.635 0.680 0.640 0.591 0.681 0.536 0.701 0.716
16×/512 0.497 1.000 0.161 0.399 0.358 0.817 0.634 0.618 0.611 0.584 0.136 0.316 0.282
16×/128 0.577 0.161 1.000 0.900 0.865 0.273 0.298 0.277 0.265 0.303 0.952 0.887 0.678
16×/256 0.705 0.399 0.900 1.000 0.956 0.527 0.572 0.495 0.410 0.542 0.920 0.936 0.744
8×/128 0.726 0.358 0.865 0.956 1.000 0.543 0.650 0.575 0.483 0.634 0.896 0.959 0.829
8×/512 0.635 0.817 0.273 0.527 0.543 1.000 0.861 0.811 0.771 0.796 0.243 0.483 0.506
8×/256 0.680 0.634 0.298 0.572 0.650 0.861 1.000 0.889 0.771 0.912 0.289 0.557 0.624
4×/256 0.640 0.618 0.277 0.495 0.575 0.811 0.889 1.000 0.873 0.941 0.240 0.532 0.600
4×/512 0.591 0.611 0.265 0.410 0.483 0.771 0.771 0.873 1.000 0.821 0.191 0.461 0.549
4×/128 0.681 0.584 0.303 0.542 0.634 0.796 0.912 0.941 0.821 1.000 0.284 0.594 0.655
10×/128 0.536 0.136 0.952 0.920 0.896 0.243 0.289 0.240 0.191 0.284 1.000 0.910 0.678
4×/64 0.701 0.316 0.887 0.936 0.959 0.483 0.557 0.532 0.461 0.594 0.910 1.000 0.814
8×/64 0.716 0.282 0.678 0.744 0.829 0.506 0.624 0.600 0.549 0.655 0.678 0.814 1.000
Appendix G Visual Sparse Steering Pseudocode
For reproducibility purposes, in Algorithm 1, we provide the pseudocode for the baselines and VS2 used in the analysis of 4.1.1.
Algorithm 1 SAE_Steering – SAE-based CLS token modification
5:     if $\text{mode}=\texttt{"reconstruction"}$" then
6:         return $\mathbf{z}_{\text{base}}$triangleright Variant: $\text{CLIP}_{\text{REC}}^{\text{F}}$
7:     else if $\text{mode}=\texttt{"amplified"}$" then
8:         return $\mathbf{z}_{\text{boost}}$triangleright Variant: $\text{CLIP}_{\text{REC}}^{\text{F}+\gamma}$
9:     else if $\text{mode}=\texttt{"steering"}$" then
10:         return $\mathbf{h}+\lambda_{\Delta z}\cdot(\mathbf{z}_{\text{boost}}-\mathbf{z}_{\text{base}})$triangleright VS2
11:     end if
12:end function
Appendix H Comparison with other Baseline Methods
We compare VS2 against SpLiCE [16], using the official implementation provided by the authors. For all three datasets, we report performance using SpLiCE with an external vocabulary of 10,000 LAION-based concepts and an $\ell_{1}$ regularization weight of 0.25, following their best reported configuration. Despite relying on no external vocabulary, VS2 consistently outperforms SpLiCE across all benchmarks highlighting the strength of sparse concept steering even in the absence of external lexical resources.
Table 8: Zero-shot top-1 accuracy (%) on CIFAR-100.
Method ViT-B/32 ViT-B/16 $\text{CLIP}_{\text{ZS}}$ 61.07 ( 0 ) 63.96 ( 0 ) $\text{SAE}_{\text{REC}}^{\text{A}}$ 58.01 ( –3.06 ) 64.05 ( +0.09 ) $\text{SAE}_{\text{REC}}^{\text{F}}$ 58.22 ( –2.85 ) 63.42 ( –0.54 ) $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ 62.69 ( +1.62 ) 66.81 ( +2.85 ) SpLiCE [16] 55.57 ( –5.50 ) 58.29 ( –5.67 ) VS2 (ours) 64.52 ( +3.45 ) 68.08* ( +4.12 )
Appendix I Effect of Top-$N$ Retrieved Neighbors
To assess the influence of retrieval size on performance, we conduct an ablation over the number of retrieved neighbors $N$ used for latent aggregation. We use a fixed ViT-B/16 (Patch-16) backbone and vary $N\in\{10,25,50,100\}$ across three datasets: CIFAR-100, CUB-200, and Tiny-ImageNet. As shown in Figure 5, both top-1 and top-5 classification accuracy steadily increase with $N$ on CIFAR-100 and Tiny-ImageNet, with diminishing returns beyond $N=50$. In contrast, performance on CUB-200 remains largely flat or slightly degrades, suggesting that retrieving more neighbors in fine-grained datasets can introduce noise due to overly similar but semantically irrelevant examples. These results indicate that the utility of neighbor-based aggregation is dataset-dependent: general object recognition tasks benefit from larger $N$, while fine-grained classification may require more careful control of retrieval scope.
Figure 5: Top-1 and Top-5 accuracy as a function of number of retrieved neighbors $N$, using ViT-B/16 (Patch-16) across datasets. Larger $N$ improves general classification but may degrade performance in fine-grained settings.
Appendix J Per-class Performance Analysis on CIFAR-100
In Table 9, we present results for the classes most affected by steering CLIP ViT-B/32 with VS2 and VS2++. We highlight both positive and negative shifts in accuracy to identify the most impacted categories. Unlike the average performance reported in the main experiments, this analysis shows that the top-10 class-level gains are generally larger in magnitude than the corresponding losses. Although the accuracy drops for misclassified categories are smaller in magnitude, they are non-negligible. As future work, it would be valuable to explore when steering should be applied or withheld to further maximize overall performance gains.
Table 9: Top-10 class gains and losses on CIFAR-100. Green = absolute gain; Red = absolute loss relative to ZS baseline.
Class ZS VS2 $\uparrow$($\Delta$) tractor $0.55$ 0.80 $\uparrow$(+0.25) forest $0.35$ 0.58 $\uparrow$(+0.23) man $0.53$ 0.74 $\uparrow$(+0.21) bus $0.51$ 0.68 $\uparrow$(+0.17) snake $0.61$ 0.76 $\uparrow$(+0.15) woman $0.57$ 0.71 $\uparrow$(+0.14) trout $0.18$ 0.32 $\uparrow$(+0.14) cockroach $0.49$ 0.62 $\uparrow$(+0.13) bridge $0.72$ 0.83 $\uparrow$(+0.11) rose $0.58$ 0.68 $\uparrow$(+0.10)
(a) VS2 – Top-10 Gains
Class ZS VS2 ($\Delta$) aquarium_fish $0.82$ 0.61 (-0.21) beetle $0.64$ 0.49 (-0.15) sweet_pepper $0.70$ 0.58 (-0.12) tulip $0.78$ 0.71 (-0.07) maple_tree $0.57$ 0.50 (-0.07) flatfish $0.55$ 0.49 (-0.06) willow_tree $0.46$ 0.41 (-0.05) lamp $0.75$ 0.70 (-0.05) lawn_mower $1.00$ 0.96 (-0.04) kangaroo $0.72$ 0.68 (-0.04)
(b) VS2 – Top-10 Losses
Class ZS VS2++ $\uparrow$($\Delta$) spider $0.48$ 0.91 $\uparrow$(+0.43) caterpillar $0.27$ 0.68 $\uparrow$(+0.41) possum $0.24$ 0.65 $\uparrow$(+0.41) tractor $0.55$ 0.96 $\uparrow$(+0.41) tiger $0.45$ 0.84 $\uparrow$(+0.39) raccoon $0.44$ 0.83 $\uparrow$(+0.39) lizard $0.36$ 0.74 $\uparrow$(+0.38) wolf $0.54$ 0.90 $\uparrow$(+0.36) shrew $0.44$ 0.79 $\uparrow$(+0.35) bear $0.51$ 0.85 $\uparrow$(+0.34)
(c) VS2++ – Top-10 Gains
Class ZS VS2++ ($\Delta$) girl $0.72$ 0.65 (-0.07) maple_tree $0.57$ 0.51 (-0.06) porcupine $0.18$ 0.13 (-0.05) ray $0.06$ 0.02 (-0.04) mouse $0.18$ 0.15 (-0.03) lawn_mower $1$ 0.98 (-0.02)
(d) VS2++ – Top-10 Losses

### v2

@@ Appendix B Sensitivity to Sparse Amplification ($\gamma$) and Steering Magnitude ($\lambda$)
Figure 4 shows that all tested combinations of $(\lambda,\gamma)$ outperform the zero-shot baseline, though to varying degrees. Each dataset exhibits a diagonal band of near-optimal settings where $\lambda\cdot\gamma\in[2,3]$ tends to yield peak accuracy. For example, CIFAR-100 peaks at $\lambda^{*}=2.1$ and $\gamma^{*}=1.5$. Beyond this band, increasing $\lambda$ or $\gamma$ causes performance to degrade likely due to over-amplification of sparse features and/or embedding distortion. The consistent contour patterns across datasets suggest that VS2 is robust to moderate variation in its hyperparameters and that good settings generalize well across domains.
@@ Appendix C Training top-$k$ Sparse Autoencoders
We follow CLIP (Radford et al., 2021b) with a ViT-B/32 and ViT-B/16 vision backbones, intercepting the output of each encoder layer for the $CLS$ token. Specifically, we train top-$k$ SAEs on the $CLS$ embeddings for each chosen layer. We use $k=64$ and $k=256$ as the maximum active features within the latent space for ViT-B/32 and ViT-B/16 respectively, and we set a “dead feature” threshold of 100 i.e., any feature seldom activated is pruned. We also use an expansion factor of 4 relative to the input embedding dimension, resulting in 3,072 latent units. Training largely follows Gao et al. (2024a) and uses  EleutherAI (2024), with a linear learning-rate schedule and warmup from $5\times 10^{-4}$.
@@
where $\mathbf{X}\in\mathbb{R}^{B\times d}$ is a batch of CLS embeddings, $\hat{\mathbf{X}}$ denotes their SAE reconstructions, and $\bar{\mathbf{X}}$ is the batch mean (so the denominator equals the total variance). FVU is the complement of the coefficient of determination ($1-R^{2}$); an FVU of 0 indicates perfect reconstruction, while an FVU of $1$ corresponds to predicting only the mean. In Table 9, we report FVU results for ViT-B/32 with $k=64$ and ViT-B/16 with $k=256$ across all three datasets.
@@ Appendix D Sensitivity in Retrieval-Augmented Generation (RAG)
When an external image corpus is available, VS2 can be extended using a Retrieval-Augmented Generation (RAG) pipeline. We use DINOv2 (Oquab et al., 2023) to retrieve top-$k$ images most similar to the input query and compute an enhanced embedding by combining the query with the retrieved set:
@@
The parameter $\alpha\in[0,1]$ controls the trade-off between using the original query and the retrieved set. We sweep over values of $\alpha$ and $k$ to assess their impact on zero-shot classification performance. Figure 5 shows results for CIFAR-100, and Tiny-ImageNet. CIFAR-100 and Tiny-ImageNet display a similar trend: larger $\alpha$ (i.e., more reliance on the query) typically degrades performance. On Tiny-ImageNet, setting $\alpha=0$, completely ignoring the input query and relying purely on retrieved features, yields the best result. Across all datasets, large $k$ eventually hurts performance, confirming that RAG benefits from focused rather than broad context. The trade-off parameter $\alpha$ and the retrieval depth $k$ play dataset-dependent roles in RAG-enhanced pipelines. For fine-grained domains, fewer, high-confidence neighbors and a low $\alpha$ might work best; for noisy domains, more reliance on the retrieval images might be more beneficial.
@@
Appendix E Steering Across Layers of the CLS Token
Our default method reconstructs and applies steering to the CLS token embedding at the final layer of the vision encoder. To study depth effects, we evaluate steering the CLS token across multiple final layers of the transformer. We use ViT-B/16 on CIFAR-100 with expansion factor 4 and 128 sparse activations. Table 10 reports top-1 accuracy when steering the last 1, 2, or 3 layers. Steering only the final layer yields the best performance. Adding earlier layers causes a progressive drop in accuracy, falling below the zero-shot baseline when steering the last three layers.
Table 10: Effect of steering depth on CLS. Top-1 accuracy (%) on CIFAR-100 (ViT-B/16) as the number of steered final layers increases.
Steered Layers Accuracy (%)
12 68.08
11 + 12 65.72
10 + 11 + 12 59.36
Applying steering at multiple layers likely introduces compounding perturbations that propagate forward, making later representations harder to align with the classifier. CLS steering is most effective at the final layer. Future work could explore coordinated multi-layer steering to avoid error accumulation.
Appendix F Ablation Study of Expansion factor and top-$k$
SAE configuration Top-1 $\uparrow$ Top-5 $\uparrow$
4 $\times$, kk=128 64.61 87.95
8 $\times$, kk=128 64.56 87.76
4 $\times$, kk=64 64.54 87.79
16 $\times$, kk=64 64.54 87.78
8 $\times$, kk=64 64.52 87.96
10 $\times$, kk=128 64.43 87.81
16 $\times$, kk=512 64.42 87.71
8 $\times$, kk=512 64.40 87.91
4 $\times$, kk=256 64.34 87.62
8 $\times$, kk=256 64.29 87.80
16 $\times$, kk=256 64.28 87.75
4 $\times$, kk=512 64.12 87.87
16 $\times$, kk=128 64.10 87.79
Appendix G Visual Sparse Steering Pseudocode
For reproducibility purposes, in Algorithm 1, we provide the pseudocode for the baselines and VS2 used in the analysis of 4.1.
Algorithm 1 SAE_Steering – SAE-based CLS token modification
5:  if $\text{mode}=\texttt{"reconstruction"}$ then
6:   return $\mathbf{z}_{\text{base}}$ $\triangleright$ Variant: $\text{CLIP}_{\text{REC}}^{\text{F}}$
7:  else if $\text{mode}=\texttt{"amplified"}$ then
8:   return $\mathbf{z}_{\text{boost}}$ $\triangleright$ Variant: $\text{CLIP}_{\text{REC}}^{\text{F}+\gamma}$
9:  else if $\text{mode}=\texttt{"steering"}$ then
10:   return $\mathbf{h}+\lambda_{\Delta z}\cdot(\mathbf{z}_{\text{boost}}-\mathbf{z}_{\text{base}})$ $\triangleright$ VS2
11:  end if
12:end function
Appendix H Comparison with other Baseline Methods
We compare VS2 against SpLiCE (Bhalla et al., 2024), using the official implementation provided by the authors. For all three datasets, we report performance using SpLiCE with an external vocabulary of 10,000 LAION-based concepts and an $\ell_{1}$ regularization weight of 0.25, following their best reported configuration. Despite relying on no external vocabulary, VS2 consistently outperforms SpLiCE across all benchmarks highlighting the strength of sparse concept steering even in the absence of external lexical resources.
Table 12: Zero-shot top-1 accuracy (%) on CIFAR-100.
Method ViT-B/32 ViT-B/16
$\text{CLIP}_{\text{ZS}}$ 61.07 ( 0 ) 63.96 ( 0 )
$\text{SAE}_{\text{REC}}^{\text{A}}$ 58.01 ( –3.06 ) 64.05 ( +0.09 )
$\text{SAE}_{\text{REC}}^{\text{F}}$ 58.22 ( –2.85 ) 63.42 ( –0.54 )
$\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ 62.69 ( +1.62 ) 66.81 ( +2.85 )
SpLiCE (Bhalla et al., 2024) 55.57 ( –5.50 ) 58.29 ( –5.67 )
VS2 (ours) 64.52 ( +3.45 ) 68.08 ( +4.12 )
Appendix I Effect of Top-$N$ Retrieved Neighbors
To assess the influence of retrieval size on performance, we conduct an ablation over the number of retrieved neighbors $N$ used for latent aggregation. We use a fixed ViT-B/16 (Patch-16) backbone and vary $N\in\{10,25,50,100\}$ across three datasets: CIFAR-100, CUB-200, and Tiny-ImageNet. As shown in Figure 6, both top-1 and top-5 classification accuracy steadily increase with $N$ on CIFAR-100 and Tiny-ImageNet, with diminishing returns beyond $N=50$. In contrast, performance on CUB-200 remains largely flat or slightly degrades, suggesting that retrieving more neighbors in fine-grained datasets can introduce noise due to overly similar but semantically irrelevant examples. These results indicate that the utility of neighbor-based aggregation is dataset-dependent: general object recognition tasks may benefit from larger $N$, while fine-grained classification may require more careful control of retrieval scope.
Figure 6: Top-1 and Top-5 accuracy as a function of number of retrieved neighbors NN, using ViT-B/16 (Patch-16) across datasets. Larger NN improves general classification but may degrade performance in fine-grained settings.
Appendix J Per-class Performance Analysis on CIFAR-100
In Table 17, we present results for the classes most affected by steering CLIP ViT-B/32 with VS2 and VS2++. We highlight both positive and negative shifts in accuracy to identify the most impacted categories. Unlike the average performance reported in the main experiments, this analysis shows that the top-5 class-level gains are generally larger in magnitude than the corresponding losses. Although the accuracy drops for misclassified categories are smaller in magnitude, they are non-negligible. As future work, it would be valuable to explore when steering should be applied or withheld to further maximize overall performance gains.
Appendix K Prototype-Aware Sparse Steering Vectors
Our core methods, VS2 and its retrieval-augmented variant VS2++, enhance zero-shot classification by steering CLIP embeddings along directions identified by a top-$k$ Sparse Autoencoder (SAE). These directions correspond to latent features that, ideally, align with class-discriminative concepts. Steering in these directions upweights what the model has learned to be important during reconstruction. This raises a central hypothesis: the reconstruction task itself is to some extent sufficient to uncover features that are also relevant for classification. In other words, there is a meaningful overlap between features that are important for reconstructing the CLS token and those that are predictive for the downstream task. In this section, drawing inspiration from prototype theory (Rosch, 1973), we investigate whether incorporating prototype information during SAE training can better align the features important for reconstruction with those that are critical for downstream classification.
CIFAR-100 CUB-200 Tiny-IN
Method ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16
Zero-shot (no retrieval)
$\text{CLIP}_{\text{ZS}}$ 61.07 ( 0 ) 63.96 ( 0 ) 51.76 ( 0 ) 55.06 ( 0 ) 56.64 ( 0 ) 61.08 ( 0 )
$\text{SAE}_{\text{REC}}^{\text{A}}$ 58.01 ( –3.06 ) 64.05 ( +0.09 ) 47.45 ( -4.31 ) 51.81 ( -3.25 ) 30.56 ( –26.08 ) 52.96 ( -8.12 )
$\text{SAE}_{\text{REC}}^{\text{F}}$ 58.22 ( –2.85 ) 63.42 ( –0.54 ) 48.08 ( -3.68 ) 51.43 ( -3.63 ) 36.33 ( –20.31 ) 54.84 ( -6.24 )
$\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ 62.69 ( +1.62 ) 66.81 ( +2.85 ) 49.41 ( -2.35 ) 53.28 ( -1.78 ) 39.49 ( –17.15 ) 58.81 ( -2.27 )
VS2 (ours) 64.52 ( +3.45 ) 68.08 ( +4.12 ) 52.69 ( +0.93 ) 56.14 ( +1.08 ) 58.14 ( +1.50 ) 62.92 ( +1.84 )
VS2 + PASS (ours) 70.64 ( +9.57 ) 68.23 ( +4.27 ) 52.97 ( +1.21 ) 56.63 ( +1.57 ) 57.94 ( +1.30 ) 62.98 ( +1.90 )
Oracle steering with known prototypes.
Following prototype theory, we collect for every class the ten images that ViT-B/32 CLIP classifies with the highest confidence; these serve as oracle prototypes. Given the true class label $y$, we build a prototype steering vector by averaging their latent sparse features. VS2 using oracle prototypes can lift CLIP to 97.5% on CIFAR-100, 91.04% on CUB-200, and 90.1% on Tiny-ImageNet. This confirms the existence of discriminative sparse directions. In Appendix L we further examine the discriminative ability (measured by orthogonality) of these steering vectors. Generally, these prototype vectors have a low cosine similarity, yet a non-negligible tail reveals strongly overlapping directions between visually or semantically close categories.
Prototype-aligned SAE (PASS).
Table 14: Trade-off between reconstruction fidelity and prototype alignment. Increasing $w_{\text{aux}}$ improves classification accuracy but degrades FVU reconstruction.
$w_{\text{aux}}$ FVU $\downarrow$ Accuracy (%) $\uparrow$
0.1 0.3437 68.80
0.5 0.4887 70.40
1.0 0.5393 70.61
2.0 0.5702 70.76
Reconstruction vs. Alignment Trade-off
Sparse Autoencoders (SAEs) trained for reconstruction can also be optimized to align their latent features with class-level prototypes. However, this introduces a trade-off between two competing objectives: fidelity of reconstruction and discriminative alignment.
To investigate this trade-off, we introduce a weighting coefficient $w_{\text{aux}}$ that controls the strength of prototype alignment relative to the reconstruction loss. As $w_{\text{aux}}$ increases, alignment is encouraged more strongly. Table 14 reports the resulting changes in reconstruction loss (measured by FVU) and top-1 classification accuracy on CIFAR-100 using ViT-B/16, with 128 sparse latents and expansion factor 4.
We observe that increasing $w_{\text{aux}}$ consistently improves classification performance, up to a point, even though it introduces more reconstruction error. This aligns with our hypothesis: while exact input reconstruction encourages general feature coverage, alignment with class prototypes promotes discriminative feature extraction. These results highlight the flexibility of SAE-based steering to balance interpretability and performance depending on downstream objectives.
Appendix L Are exemplar-derived directions really distinct?
@@ Appendix I Effect of Top-$N$ Retrieved Neighbors
A desirable property of class–specific steering vectors is orthogonality: pushing an embedding toward class A should not simultaneously raise its score for class B. Using the oracle prototypes as described in Appendix K, we compute for every CIFAR-100 class a prototype steering vector and measure the pair-wise cosine similarity at layer 11 of the ViT-B/32 encoder. Most pairs have low similarity (mean=0.23), yet a non-negligible tail reveals strongly overlapping directions. Table 15 lists the ten highest-overlap pairs.
@@
Table 16: Training-time compute and parameter cost per sample for various adaptation strategies on the CLIP ViT-B/32 vision encoder.
Method Train FLOPs / sample (fwd+bwd) GFLOPs Trainable Params
SAE (cached activations) 14,164,992 0.014 4,722,432
SAE (non-cached; incl. extraction) 8,743,645,440 8.744 4,722,432
LoRA (rank = 16) 26,365,388,544 26.365 589,824
Full fine-tuning (vision encoder) 26,188,441,344 26.188 87,456,000
@@
Table 17: Top-5 class gains and losses on CIFAR-100. Green = absolute gain; Red = absolute loss relative to ZS baseline.
Class ZS VS2 $\uparrow$($\Delta$) tractor $0.55$ 0.80 $\uparrow$(+0.25) forest $0.35$ 0.58 $\uparrow$(+0.23) man $0.53$ 0.74 $\uparrow$(+0.21) bus $0.51$ 0.68 $\uparrow$(+0.17) snake $0.61$ 0.76 $\uparrow$(+0.15)
(a) VS2 – Top-5 Gains
Class ZS VS2 ($\Delta$) aquarium_fish $0.82$ 0.61 (-0.21) beetle $0.64$ 0.49 (-0.15) sweet_pepper $0.70$ 0.58 (-0.12) tulip $0.78$ 0.71 (-0.07) maple_tree $0.57$ 0.50 (-0.07)
(b) VS2 – Top-5 Losses
Class ZS VS2++ $\uparrow$($\Delta$) spider $0.48$ 0.91 $\uparrow$(+0.43) caterpillar $0.27$ 0.68 $\uparrow$(+0.41) possum $0.24$ 0.65 $\uparrow$(+0.41) tractor $0.55$ 0.96 $\uparrow$(+0.41) tiger $0.45$ 0.84 $\uparrow$(+0.39)
(c) VS2++ – Top-5 Gains
Class ZS VS2++ ($\Delta$) girl $0.72$ 0.65 (-0.07) maple_tree $0.57$ 0.51 (-0.06) porcupine $0.18$ 0.13 (-0.05) ray $0.06$ 0.02 (-0.04) mouse $0.18$ 0.15 (-0.03)
(d) VS2++ – Top-5 Losses
Appendix M Computational Overhead: Additional Details
Training-time FLOPs.
To quantify the cost of training the SAE versus alternative adaptation methods, Table 16 reports per-sample training FLOPs (forward + backward) and the number of trainable parameters for the CLIP ViT-B/32 vision encoder. Even when activation extraction is included, SAE training remains more than $3\times$ cheaper than LoRA or full fine-tuning. With cached activations, as in our VS2 setup, the cost drops to just 0.014 GFLOPs per sample.
TPT FLOP accounting.
For completeness, we outline how the $16.6\times$ test-time overhead factor for TPT (Shu et al., 2022) is obtained. A single CLIP ViT-B/32 forward pass on a $224\times 224$ image with 100 text prompts costs $8.7$ GFLOPs for the vision encoder and $581.8$ GFLOPs for the text encoder, for a total of approximately $0.59$ TFLOPs. In our implementation, TPT performs $63$ augmented views per image and runs $4$ optimization steps; each step requires a full CLIP forward pass over all views and a backward pass through the text encoder. This yields about $9.217$ TFLOPs of adaptation compute per test image, followed by one final forward-only evaluation of $0.591$ TFLOPs, for a total of roughly $9.807$ TFLOPs. Dividing by the $0.59$ TFLOPs of plain CLIP gives the reported $\approx 16.6\times$ test-time compute overhead.
@@
SAE cfg. 16×/64 16×/512 16×/128 16×/256 8×/128 8×/512 8×/256 4×/256 4×/512 4×/128 10×/128 4×/64 8×/64 16×/64 1.000 0.497 0.577 0.705 0.726 0.635 0.680 0.640 0.591 0.681 0.536 0.701 0.716 16×/512 0.497 1.000 0.161 0.399 0.358 0.817 0.634 0.618 0.611 0.584 0.136 0.316 0.282 16×/128 0.577 0.161 1.000 0.900 0.865 0.273 0.298 0.277 0.265 0.303 0.952 0.887 0.678 16×/256 0.705 0.399 0.900 1.000 0.956 0.527 0.572 0.495 0.410 0.542 0.920 0.936 0.744 8×/128 0.726 0.358 0.865 0.956 1.000 0.543 0.650 0.575 0.483 0.634 0.896 0.959 0.829 8×/512 0.635 0.817 0.273 0.527 0.543 1.000 0.861 0.811 0.771 0.796 0.243 0.483 0.506 8×/256 0.680 0.634 0.298 0.572 0.650 0.861 1.000 0.889 0.771 0.912 0.289 0.557 0.624 4×/256 0.640 0.618 0.277 0.495 0.575 0.811 0.889 1.000 0.873 0.941 0.240 0.532 0.600 4×/512 0.591 0.611 0.265 0.410 0.483 0.771 0.771 0.873 1.000 0.821 0.191 0.461 0.549 4×/128 0.681 0.584 0.303 0.542 0.634 0.796 0.912 0.941 0.821 1.000 0.284 0.594 0.655 10×/128 0.536 0.136 0.952 0.920 0.896 0.243 0.289 0.240 0.191 0.284 1.000 0.910 0.678 4×/64 0.701 0.316 0.887 0.936 0.959 0.483 0.557 0.532 0.461 0.594 0.910 1.000 0.814 8×/64 0.716 0.282 0.678 0.744 0.829 0.506 0.624 0.600 0.549 0.655 0.678 0.814 1.000